An ice cream store offers 30 flavors of ice cream. How many different triple-decker cones are possible? (Note: Assume that different orders of the same flavors are not different cones. Thus a scoop of rocky road followed by two scoops of mint chocolate is the same as one scoop of mint chocolate followed by one scoop of rocky road followed by a second scoop of mint chocolate.)
Answer by stanbon(75887)

An ice cream store offers 30 flavors of ice cream. How many different triple-decker cones are possible?
----
Since order makes no difference
the answer is a combination.
Ans: 30C3 = (30*29*28)/*(1*2*3) = = 5*29*28 = 4060
----------
Cheers,
Stan H.
============